The population growth of an animal species is described by F(t)=500+90 log3 (2t+1) where t is measured in months. Find the population of this species in an area 40 month(s) after the species is introduced.
Answer by oscargut(686)

F(40)=500+90 log3 (2(40)+1)=500+90log3 (81)=500+90(4)=860
Answer: 860